The following is the mass spectrum of Niacin. Can you please interpret each of the peaks seen in the mass spectrum and what they correspond to on the Niacin compound?Niacin Mass Spectrum Relative Intensity LES .45 1718 2025262728 No 89 9495 III. 6162366566768 70 60 72757 11,80 80 122| 20 30
Answer #1

Here I have drawn only intense peak .m-18 +Ho -Hao Niacin M/Z=105 M*+ miz- 123 M-45 + Mot ml2- 123 oH m-46 m-46 M/z -123 H-coo m2= 77 OH m-45 m-26 m/2 78 CEC-H +miz = 123 fs the molecular fon peak comesponds 123-105 18 ie. Hao doss Hao to peak ml 105 123 78 45 thrs fndicates the OF ROH